# Large Language Models and Code Security: A Systematic Literature Review

### v1

@@ VII-B LLMs for Detecting and Fixing Vulnerabilities in Self-produced Code
Code generated by LLMs can contain vulnerabilities, as we have previously discussed in Section IV. However, researchers have shown that it is possible to re-prompt LLMs to identify and fix those vulnerabilities. By doing that, they have explored the self-evaluation approach and they both concluded that the code can be significantly improved through iterative refinement.
@@ VII-C Fine-tuning for Vulnerability Detection and Fixing
Moreover, Zhang et al. [119] fine-tuned LLMs on GitHub workflow tasks. Their work explored LLMs effectiveness in generating workflows, identifying defects, and repairing them, focusing on syntactic errors and code injection vulnerabilities. Although GPT-3.5 often produced expected workflows, it occasionally generated invalid or defective ones. In contrast, CodeLlama was less likely to generate expected workflows, but had a higher probability of producing valid workflows. In terms of error detection, off-the-shelf StarChat achieved the best results in detecting syntactic errors, with a 100% F1 score, while fine-tuned GPT-3.5 performed best at detecting code injection vulnerabilities, with about 99% F1 score. However, LLMs can create workflows with code injection vulnerabilities, so it is important to be cautious when using workflows generated by LLMs. Last, LLMs excel at fixing syntactic errors, but they struggle when it comes to fixing code injection vulnerabilities. Fine-tuned CodeLlama and StarChat performed worse on new, but similar workflow tasks. In conclusion, LLMs struggled when generating large and valid workflows, and seemed ineffective at fixing code injection vulnerabilities.
@@ VIII Prompting strategies for detecting and fixing vulnerabilities
The effectiveness of LLMs for code vulnerability detection and fixing can vary significantly, depending on how the prompts are structured. Well-crafted prompts can guide LLMs towards more accurate and efficient solutions, while incomplete or ambiguous prompts may result in partial or ineffective solutions [46]. This section explores various prompt strategies to enhance the performance of LLMs in such tasks. To summarize the studies for the reader, Table V lists the different prompting strategies, and corresponding studies, discussed in this section.
TABLE V: Prompt engineering techniques used in the studies for vulnerability detection and fixing.
@@ VIII-B Few-shot Prompting
While zero-shot prompting involves providing a model with only a general instruction or task description, few-shot prompting is a technique where the model is given a few task-specific examples, before performing the task [128]. An example of a prompt using the few-shot technique is shown in Figure 5, where a set of task-specific examples are provided to the model, as part of the input.
Example 1: <code> - Code is vulnerable to arbitrary file access. Example 2: <code> - Code is vulnerable to command injection. Now, analyze the following code for vulnerabilities: <code>
@@ VIII-C Chain-of-Thought Prompting
Example 1: <code> - Takes user input as filename. - Opens the file specified by user input. - What could go wrong? If the user enters a malicious filename or path, it could point to sensitive files. - Conclusion: code is vulnerable to arbitrary file access. Example 2: <code> - Takes user input as a shell command. - Executes the command directly. - What could go wrong? The user could inject additional commands, leading to unintended execution on the system. - Conclusion: code is vulnerable to command injection. Now, analyze the following code for vulnerabilities: <code>
Figure 6: Example of a prompt, using the CoT prompting technique. In the gray boxes, the code examples and the reasoning steps provided to the LLM. In green, the main body of the instructions. In blue, the code to be analyzed.
@@
Experimental findings from Liu [61], indicate that CoT enhances vulnerability detection, while few-shot prompting negatively impacts it. Similarly, [118] proposed to incorporate CoT prompting within their framework for detecting and explaining vulnerabilities, showing that this approach can improve LLMs ability to identify and clarify security issues.
@@ VIII-D In-context Prompting
Steenhoek et al. [56] experimented two different in-context approaches. In the first approach, they provided the model with only in-context examples of vulnerable and safe code, achieving a balanced accuracy of about 58%. In the second approach, they combined in-context prompting with CoT in a two-step process. They started by presenting the model with a few examples of vulnerable and safe code (in-context), following with examples of step-by-step reasoning (CoT). Albeit both approaches worked, the second approach resulted in a slightly lower balanced accuracy, compared to the first.
@@ VIII-E Task-oriented Prompting
Task-oriented prompting involves providing clear, specific, and detailed instructions to guide a model in completing a task [59]. Unlike in-context prompting, which uses examples within the prompt to illustrate a task, task-oriented prompting focuses on specifying We provide an example of task-oriented prompting in Figure 8, demonstrating how the user provides specific tasks that the LLM should complete.
@@ IX-B Impact of Poisoned Training Data on Ability to Detect and Fix Vulnerabilities
Although research on data poisoning in LLMs exists, studies specifically examining its effects on vulnerability detection and fixing are currently absent. This gap in the literature will be further discussed in the Open challenges, in Section X-B.
@@ X Discussion
In Secton III, as a part od our methodology, we have defined three RQs that steered the focus of this study. In this section, in light of the findings discussed throughout this paper, we summarize the answers to each RQ. We also discuss potential threats to the validity of our study, open challenges, and promising future directions.
@@ X-B Open Challenges
Despite the advancements in understanding poisoning attacks on LLMs, there remains a gap in research that specifically addresses how poisoning attacks affect the ability of LLMs to detect and fix vulnerabilities. Given that poisoned models have been demonstrated to suggest insecure code, it is reasonable to hypothesize that poisoning could have affected the LLM ability to effectively detect and fix security vulnerabilities in code. This lack of research presents an open challenge for the field.
In addition to this research gap, there are several areas for improvement of LLMs. There is a need to minimize the number of vulnerabilities present in the code generated by these models. Also, the number of false positives when LLMs attempt to detect vulnerabilities in code should be reduced. Another area of improvement is to enhance the LLMs ability to detect vulnerabilities in obfuscated code. Furthermore, LLMs need improvements to be capable of detecting and fixing vulnerabilities in real-world scenarios, where code complexity and variations make it more challenging to identify and fix security issues compared to controlled or simplified environments. Finally, research must focus on reducing the impact of poisoned training data on the security of the code generated by LLMs.

### v2

@@ VII-B LLMs for Detecting and Fixing Vulnerabilities in Self-produced Code
Code generated by LLMs can contain vulnerabilities, as we have previously discussed in Section IV. However, researchers have shown that it is possible to re-prompt LLMs to identify and fix those vulnerabilities. By doing that, they have explored the self-evaluation approach and concluded that the code can be significantly improved through iterative refinement.
@@ VII-C Fine-tuning for Vulnerability Detection and Fixing
Zhang et al. [107] fine-tuned LLMs on GitHub workflow tasks. They explored LLMs effectiveness in generating workflows, identifying defects, and repairing them, focusing on syntactic errors and code injection vulnerabilities. Although GPT-3.5 produced expected workflows, it occasionally generated defective ones. In contrast, Code Llama was less likely to generate expected workflows but had a higher probability of producing valid workflows. In terms of error detection, off-the-shelf StarChat achieved the best results in detecting syntactic errors, with a 100% F1 score, while fine-tuned GPT-3.5 performed best at detecting code injection vulnerabilities, with about 99% F1 score. However, LLMs can create workflows with code injection vulnerabilities, so it is important to be cautious when using workflows generated by LLMs.
@@ VIII Prompting strategies for detecting and fixing vulnerabilities
The effectiveness of LLMs for code vulnerability detection and fixing can vary significantly, depending on how the prompts are structured. Well-crafted prompts can guide LLMs towards more accurate and efficient solutions, while incomplete or ambiguous prompts may result in partial or ineffective solutions [46]. This section explores the impact of various prompting strategies on the detection and fixing performance of LLMs. To summarize the studies for the reader, Table VII lists the different prompting strategies, and corresponding studies.
TABLE VII: Prompt engineering techniques used in the studies for vulnerability detection and fixing.
@@ VIII-B Few-shot Prompting
While zero-shot prompting involves providing a model with only a general instruction or task description, few-shot prompting is a technique where the model is typically given a few task-specific examples, before performing the task [116], with n-shot prompting referring to cases that use a specific number n of examples. An example of a prompt using the few-shot technique is shown in Figure 5, where a set of task-specific examples are provided to the model, as part of the input.
Example 1: <code1> - Code is vulnerable to arbitrary file access. Example 2: <code2> - Code is vulnerable to command injection. Now, analyze the following code for vulnerabilities: <code3>
@@ VIII-C Chain-of-Thought Prompting
Example 1: <code1> - Takes user input as filename. - Opens the file specified by user input. - What could go wrong? If the user enters a malicious filename or path, it could point to sensitive files. - Conclusion: code is vulnerable to arbitrary file access. Now, analyze the following code for vulnerabilities: <code2>
Figure 6: Example of a prompt, using the CoT prompting technique. In the gray boxes, the code example and the reasoning steps provided to the LLM. In green, the main body of the instructions. In blue, the code to be analyzed.
@@
Experimental findings from Liu et al. [57], indicate that CoT enhances vulnerability detection, while few-shot prompting negatively impacts it. Similarly, Mao et al. [106] proposed to incorporate CoT prompting within their framework for detecting and explaining vulnerabilities, showing that this approach can improve LLMs ability to identify and clarify security issues.
@@
In conclusion, CoT prompting outperforms zero-shot and few-shot prompting techniques in vulnerability detection and fixing. The literature offers various advanced methods and frameworks to further enhance CoT capabilities, making it even more promising than it stands. However, some of these approaches, such as ToT, often face computational challenges.
@@ VIII-D In-context Prompting
Moreover, Steenhoek et al. [52] experimented two different in-context approaches. In the first approach, they provided the model with only in-context examples of vulnerable and safe code, achieving a balanced accuracy of about 58%. In the second approach, they combined in-context prompting with CoT in a two-step process. They started by presenting the model with a few examples of vulnerable and safe code (in-context), following with examples of step-by-step reasoning (CoT). Albeit both approaches worked, the second approach resulted in a slightly lower balanced accuracy, compared to the first.
@@
The findings from these studies suggest that in-context prompting, compared to context-free approaches, significantly enhances vulnerability detection and fixing, particularly when the provided context is carefully tailored to the task.
@@ VIII-E Task-oriented Prompting
Task-oriented prompting involves providing clear, specific, and detailed instructions to guide a model in completing a task [55]. Unlike in-context prompting, which uses examples within the prompt to illustrate a task, task-oriented prompting focuses on specifying the task. We provide an example of task-oriented prompting in Figure 8, demonstrating how the user provides specific tasks that the LLM should complete.
@@
Overall, task-oriented prompts improve vulnerability detection by considerable margins. Encouraging prompts also show promise, though the accuracy of their results decreases with larger codebases.
@@ IX-B Impact of Poisoned Training Data on Ability to Detect and Fix Vulnerabilities
Although research on data poisoning in LLMs exists, there are currently no studies that examine its particular effects on vulnerability detection and fixing tasks. This gap in the literature will be further discussed in the Open challenges, in Section X-B.
@@ X Discussion
In Section III, as a part of our methodology, we have defined three RQs that steered the focus of this study. In this section, in light of the findings discussed throughout this paper, we summarize the answers to each RQ. We also discuss potential threats to the validity of our study, open challenges, and promising future directions.
@@ X-B Open Challenges
Although LLMs show great potential in code generation and security-related tasks, they still face challenges that need to be addressed. As discussed in Section IV, LLM-generated code often introduces security vulnerabilities. Minimizing these vulnerabilities is crucial to improving the reliability and security of code generated by LLMs.
Besides the aforementioned challenges faced in producing secure code, LLMs encounter challenges also when analyzing code for detecting vulnerabilities in it. As we have uncovered in Subsection V-B, the first and main issue is the one of false positives, which severely limits current LLMs capabilities in vulnerability detection. Reducing false positive rates is necessary for ensuring higher reliability of LLMs and their broader adoption for undertaking these tasks. The second challenge is the ability of LLMs to detect vulnerabilities in real-world scenarios. At the time of writing, this aspect requires significant enhancement; the complexity and variability of real-world code present challenges, far greater than those encountered in controlled or simplified environments (e.g., most vulnerability datasets).
Third, LLMs encounter substantial challenges when it comes to fixing vulnerabilities in code. As described in Section VI, the ability of LLMs to fix complex security vulnerabilities often requires a deep contextual understanding of the code and its dependencies, deeper than what current LLMs offer. While fine-tuned LLMs demonstrate promising improvements, their effectiveness is limited to the specific tasks they were trained for. Therefore, as it stands, LLMs need to be significantly improved for allowing them to be used as standalone tools for fixing vulnerabilities.
Finally, as discussed in Subsection IX-A, the literature reveals that poisoned data can influence LLMs to generate vulnerable code. To address this challenge, future research should focus on mitigating the potential impact that malicious data could have on the security of LLM-generated code. Despite the advancements in understanding poisoning attacks on LLMs, as discussed in Subsection IX-B, there is a lack of literature work that addresses how poisoning attacks could also affect the ability of LLMs to detect and fix vulnerabilities. Given that poisoned models have been shown being capable of suggesting insecure code, it is reasonable to hypothesize that poisoning could affect LLMs ability to effectively detect and fix security vulnerabilities in code (whether computer- or human-generated).
@@ XI Conclusion
Acknowledgment
This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. We would also like to express our gratitude to Markus Kuchler for his continuous support throughout the course of our research. His participation has been invaluable to the progress of this work.